Hello,
I was wondering how you write the equation of a line that is drawn on a graph. I think you would use y=mx+b, but I am not sure how you would plug the numbers in. The problem just shows a graph with a diagonal line and the y intercept is 0,2. The x intercept is 2,0. I know this question may sound confusing because I can't draw the graph but if you could help that would be great!
Thanks




























Answer by stanbon(75887)

y=mx+b, the y intercept is 0,2. The x intercept is 2,0.

--------------
You need to find the slope m, and the value of b.
m=[2-0]/[0-2}=2/-2=-1
Use the point (0,2) and the value of m in the equation
to find b as follows:
2=-1(0)+b
b=2
So the equation is y=-x+2
Cheers,
Stan H.